# A Fast and Model Based Approach for Evaluating Task-Competence of Antagonistic Continuum Arms

### v1

@@ IV-D Evaluation of Task Shapes
Although so far we have considered design problems with task-shape requirements that specify the shape of the entire arm, in reality many tasks only place requirements on the pose of the arm’s tip. While there are an infinite number of curves that satisfy the geometry of such tasks, we currently lack methods to identify the smaller set of shapes that an arm can feasibly attain. While this challenge deserves comprehensive consideration in future work, here we demonstrate one approach using our wrench-hull methods to compare the attainability of different arm shapes for reaching the same tip pose under the same load (Figure 7). Three different task-shapes for reaching the same tip-pose are shown in Fig. 7 - see our SI how we generated them. For load, we consider a tip-load of $\begin{bmatrix}7&0&0\end{bmatrix}^{\textrm{T}}$, equivalent to opening a heavy drawer.
@@ V CONCLUSION
In this work, we motivated and defined the specific problem of analyzing whether a soft robot arm can accomplish desired tasks. We introduced a novel method to determine a task’s attainability, and in the event that a task is not attainable, we proposed new metrics to quantify the unattainability of the task. We demonstrated our method’s speed and interpretability on a range of common arm design problems, and yielded concrete design insights by applying the method specifically to analyze antagonistic soft robot arms.
Further work is needed to study how attainability changes over different arm dimensions, and to use these methods to inform the design of a soft robot arm to complete currently challenging tasks. In particular, it may be interesting to more carefully explore selection of target shapes for efficient movement, as well as increased pressure segmentation.

### v3

@@ IV-D Evaluation of Task Shapes
Although we have only considered design problems with requirements on the shape of the entire arm, real tasks often only place requirements on the arm’s tip’s pose. There are an infinite number of curves that satisfy the geometry of such tasks, but we currently lack methods to identify the smaller set of shapes that an arm can feasibly attain. This challenge deserves comprehensive consideration in future work - here we demonstrate how our wrench-hull methods to compare different arm shapes for reaching the same tip pose under the same load (Figure 7). Three different task-shapes for reaching the same tip-pose are shown in Fig. 7 - see our SI for details. We consider a tip-load of $\begin{bmatrix}7&0&0\end{bmatrix}^{\textrm{T}}$, equivalent to opening a heavy drawer.
@@ V CONCLUSION
In this work, we motivated and defined the problem of determining if a soft robot arm can accomplish desired tasks. We introduced a novel method to determine a task’s attainability, and when a task is not attainable, we proposed new metrics to quantify the unattainability. We demonstrated our method’s speed and interpretability across common arm design problems, and yielded concrete design insights by applying the method to antagonistic soft robot arms.
Although our work accelerates a key bottleneck that prevents researchers from designing soft robot arms to complete currently challenging tasks, future work is needed to develop requirement-driven-design methodologies that capitalize on the speed of our method. We could also explore selecting target shapes for efficient movement, or formalize the effects of increased pressure segmentation. Finally, while the authors acknowledge that extending this method into dynamic systems is necessary future work, there are also still many interesting results to explore in analyzing static requirements.